A man age is 4 times the combine age of his 2 son one of whom is 3 times as old as the other in 24 years their combine ages will be in 12 years less than their father age. what are their age.
Answer by mananth(16946)

let one son's age =x
the other son's age =3x
...
father's age = 4(x+3x)
=16x
..
after 24 years
son x+24
other son 3x+24
Father = 16x+24
...
x+24+3x+24=16x+24-12
4x+48=16x+12
16x-4x=48-12
12x=36
/12
x=3 years one son's age
second son's age = 3x= 9 years
father's age = 16x = 48 years
...
m.ananth@hotmail.ca